OpenStudy (anonymous):

Determine the number of solutions the system of 4(x + 2) = y + x and y = 2x + 6. A) No solution B) One solution C) Infinite solutions D) Impossible to determine

OpenStudy (anonymous):

you can do it a little matrix and solve

OpenStudy (anonymous):

explain plz

OpenStudy (anonymous):

4x+8=y+x 3x-y=-8 -2x+y=6 add the 2nd equation to the 1st. so x=-2 if x=-2 then y=2 so one solution

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

it makes sense because both are equations of a line. as long as they are not parallel and they are linear (meaning exponents to the power of 1 or 0) then there is only one point at which they intersect.